Q=200-2P+4i and Q=-50+3P


1
Expert's answer
2021-08-25T17:07:08-0400

If Q=200-2P+4i and Q=-50+3P, then in equilibrium:

"200-2P+4i =-50+3P,"

5P = 250 + 4i,

P = 50 + 0.8i.